our answer is incorrect. Try again.

The inventory of Flint Company on December 31, 2017, consists of the following items.

Part

Quantity

Cost per Unit

Net Realizable Value

110 600 $98 $103
111 920 62 54
112 490 82 78
113 220 175 185
120 360 211 214
121 a 1,500 16 1
122 320 247 242


a Part No. 121 is obsolete and has a realizable value of $1 each as scrap.

(a) Determine the inventory as of December 31, 2017, by the LCNRV method, applying this method to each item.

Inventory as of December 31, 2017 $Entry field with incorrect answer
(b) Determine the inventory by the LCNRV method, applying the method to the total of the inventory.
Inventory as of December 31, 2017 $

Entry field with incorrect answer
Answer #1

Answer-a)- Inventory as of December 31,2020, by the LCNRV method to each item = $340100.

Explanation-

Calculation of closing inventory as of December 31,2017 for each item
By using Lower of cost or Net Realizable Value Method Closing Inventory
Part Quantity (a) Cost per unit Net Realizable value Valued on $ Per unit (b) Amount $ (C=a*b)
$ $
110 600 98 103 Cost 98 58800
111 920 62 54 Net Realizable value 54 49680
112 490 82 78 Net Realizable value 78 38220
113 220 175 185 Cost 175 38500
120 360 211 214 Cost 211 75960
121a 1500 16 1 Net Realizable value 1 1500
122 320 247 242 Net Realizable value 242 77440
Totals 340100

b)- Inventory as of December 31,2020, by the LCNRV method to the total of the inventory = $346380.

Explanation-

Calculation of closing inventory as of December 31,2017 to total of the Inventory
By using Lower of cost or Net Realizable Value Method
Part Quantity (a) Cost per unit (b) Net Realizable value (c) Total cost( d=a*b) Total NRV (e=a*c)
$ $ $ $
110 600 98 103 58800 61800
111 920 62 54 57040 49680
112 490 82 78 40180 38220
113 220 175 185 38500 40700
120 360 211 214 75960 77040
121a 1500 16 1 24000 1500
122 320 247 242 79040 77440
Totals 373520 346380